A number is divided into two parts, such that one part is 12 more than the other. If the two parts are in the ratio 4 : 3, find the number and the two parts.

Found 3 solutions by Theo, ikleyn, greenestamps:
Answer by Theo(13342)

let one part be x + 12
let the other part be x.
your ratio becomes:
(x + 12) / x = 4/3
cross multiply to get:
4x = 3 * (x + 12)
simplify to get:
4x = 3x + 36
subtract 3x from both ides of the equation to get:
x = 36

your solution is:
the number is equal to 36
the first part is 48.
the second part is 36.

confirm your solution is correct by doing the following:
the ratio is 48/36 = 4/3
cross multiply to get:
4 * 36 = 3 * 48
simplify to get:
144 = 144
this confirms the solution is correct.

Answer by ikleyn(52788)

.

The correct answer is THIS:

        The number is 84 and the parts are 48 and 36.



Answer by greenestamps(13200)



The two parts are in the ratio 4:3, so represent the two numbers with 4x and 3x.

The difference between the two numbers is 12:

4x-3x=12
x=12

Then the two numbers are 4x=48 and 3x=36; the sum is 84.

ANSWERS: 84, 48, and 36